OpenStudy (anonymous):

Solve........ -15(4+3x)-8x=26-6(x-9)

OpenStudy (gw2011):

Multiplying the equation you get: -60-45x-8x=26-6x+54 Combine like terms and you get: -60-53x=-6x+80 Again, combine like terms and you get: -60-80=53x-6x Rearranging this equation you get: 53x-6x=-60-80 47x=-140 x=-140/47=-2.979